# Index Terms: Machine-generated code detection, large language models, code generation, empirical analysis

### 24 Jan 2024 v2

@@ Abstract
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine and human-authored code. Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct stylized patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
Index Terms:
Machine-generated code detection, large language models, code generation, empirical analysis
1 Introduction
The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect. LLMs trained on code [3, 4, 5, 6], sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation [7, 8]. Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
@@ 1 Introduction
While capable of generating human-like programs, LLMs bring ambiguity of whether a software artifact is created by a human or machine, causing integrity and authenticity issues in software development. This indistinction can lead to various challenges, such as the misattribution of code ownership for bug/fault triage and potential vulnerabilities in machine-generated code that may go unnoticed due to overreliance on its perceived robustness. Furthermore, the ease of generating code could lead to inflated assessments of project workloads, skewing economic valuations and potentially compromising the reliability of software. This blending of human and machine efforts not only raises questions about the trustworthiness of the software but also threatens the integrity of coding as a discipline, wherein the true authorship and the effort invested in creating software artifacts become obscured. Addressing these concerns is pivotal in maintaining a transparent and secure software development lifecycle.
Recently, there has been a growing research trend in detecting machine-generated texts [9, 10]. Perturbation-based methods like DetectGPT [11], have achieved state-of-the-art results in identifying machine-generated text. DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy. And the perturbation process is performed by a pretrained language model (e.g., T5 [12]) to mask portions of the original text and then reconstruct it. While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid syntactic constraints. Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code [13].
@@
In this paper, we conduct a comparative analysis of the patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness. Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespace tokens.
Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code. Our DetectCodeGPT improves DetectGPT by capturing the distinct styles of code. Diverging from conventional techniques that depend on external pretrained model (e.g. CodeT5 [14]) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code by strategically inserting spaces and newlines. Eschewing the need for an LLM for perturbations, our method provides much higher efficiency. Since the proposed perturbation strategy well captures the stylistic distinctions between machine- and human-authored code, our DetectCodeGPT also ensures higher efficacy.
To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models, ranging from 1.3 to 7 billion parameters, and operating at two different decoding conditions. The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward [9, 15], detecting machine-authored code presents a unique challenge. Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement.
@@
  • To our knowledge, we are the first to conduct a comprehensive and thorough analysis of the differences between machine- and human-authored code. Our study sheds light on essential insights that further advance the utility of LLMs in coding.
  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct stylistic patterns of code.
@@
2 Background
2.1 Large Language Models for Code
GPT language models [16, 17] based on Transformer [18] decoder has achieved remarkable success in natural language processing tasks [19]. In the file of code generation, Codex [1] and AlphaCode [20] are pioneering works to train large language models on code. These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens [16]. Mathematically, this is captured as:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$ (1)
where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$. And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub [21, 22, 23].
Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle [14, 24, 4, 5] and also instruction fine-tuning [25, 6]. And recent large language models pretrained on a mixture of code and natural language like ChatGPT [26], GPT4 [27] and LLaMA [28] has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
2.2 Perturbation-Based Detection of Machine-Generated Text
@@ 3.1.1 Lexical Diversity
We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human lexical usage.
@@
3.1.2 Conciseness
@@ 3.1.3 Naturalness
3.2 Experimental Setup
3.3 Dataset Preparation
To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus [21], a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase. We used function signatures and their accompanying comments as prompts for the model as in [1], and we collected the corresponding bodies of these functions to represent human-written code.
While acknowledging that current models, including CodeLlama and even ChatGPT [43], may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet [7, 43], the choice of this dataset is deliberate and insightful. Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [44] dataset, offers a more representative assessment. It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
3.4 Results and Analysis
We present the results and analysis regarding each code attribute introduced in Section 3.1.
3.4.1 Token Frequency
@@
31–40 try * os len format get and True value isinstance else TypeError str ‘ __init__ is ¿ // the
@@
Finding 1: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
3.4.2 Syntax Element Distribution
Figure 1: Syntax element distribution of the studied code corpus
The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machines favor reusing established identifiers over creating new ones, possibly a reflection of its programming paradigm influenced by its training data.
Literal: Machine-authored code consistently shows a marginally higher inclination towards literals. Across both temperatures, the machine exhibits an increase in the literal proportion as opposed to the human-written code. This indicates a potential proclivity of machine-authored code to process raw data more frequently. This could be attributed to the machine’s broader exposure to varied tasks involving data manipulations.
@@ 3.4.2 Syntax Element Distribution
Finding 2: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
3.4.3 Zipf’s and Heaps’ Laws
@@
3.4.4 Number of Tokens and Lines
@@
Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
3.4.5 Token Likelihood and Rank
Figure 4: Distribution of naturalness scores for machine- and human-authored code
@@ 3.4.5 Token Likelihood and Rank
4 Detecting Machine Generated Code
The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespaces. This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference. Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code. Our approach adapts the perturbation-based detection framework DetectGPT [11] to the code domain.
4.1 Problem Formulation
We formulate machine-generated code detection as a classification task that predicts whether a given code snippet $x$ is produced by a source model $p_{\theta}$. For this purpose, we transform $x$ to an equivalent form $\tilde{x}$ through a perturbation process $q(\cdot|x)$. We anticipate a sharper decline in its naturalness score if $x$ is written by an LLM. The key problems here are how to define the naturalness score and how to design the perturbation process. We introduce the naturalness score and the perturbation strategy $q(\cdot|x)$ in our approach in the following sections.
4.2 Measuring Naturalness
@@ 2       $p\sim\mathcal{U}(0,1)$;
4             Let $C$ represent the set of all possible locations to insert spaces in $x$;
@@ 22      Store the perturbed code as $\tilde{x}_{i}$;
25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
@@
4.3 Perturbation Strategy
Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style. Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below. Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
4.3.1 Space Insertion
$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (4)
4.3.2 Newline Insertion
We split the codes into lines and obtain a set $L$ of lines. A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations. For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$. Formally, this is expressed as:
$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (5)
We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations. Through this step, we instill randomness at a granular stylistic level, thereby amplifying the perturbation’s efficacy. Our perturbation strategy introduces several distinct advantages over the conventional methods [11, 45] using MLM to perturb the code, which will be discussed in Section 6.2.
@@ 5 Evaluation
  • RQ4: How resilient is our method to cross-model code source detection?
5.1 Datasets
Apart from CodeSearchNet [21] described in Section 3.3, we select Python code from The Stack [22] as another evaluation dataset. Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version [46] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1]. For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation. The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
5.2 Evaluation Metric
Following prior works [11, 45], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code. Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:

### 24 Mar 2024 v3

# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

@@ Abstract
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine- and human-authored code. Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose DetectCodeGPT, a novel method for detecting machine-generated code, which improves DetectGPT by capturing the distinct stylized patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code  [^1] Our code and data are available at https://github.com/YerbaPage/DetectCodeGPT.
I Introduction
The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering tasks such as code generation. Through extensive training on ultra-large code corpora [3, 4, 5, 6], LLMs acquire the ability to generate syntactically and functionally correct code, bringing a new era of efficiency and innovation in software creation, maintenance, and evolution.
@@ I Introduction
While capable of generating human-like code, LLMs bring ambiguity of whether a software artifact is created by human or machine, causing integrity and authenticity issues in software development. This indistinction can lead to various challenges, such as misattribution of code ownership for bug triage and inflated assessments of developer workloads. Potential vulnerabilities in machine-generated code may go unnoticed due to the overreliance on its perceived robustness. The blending of human and machine efforts not only raises questions about the trustworthiness of the software but also threatens the integrity of development process, wherein the true authorship and the effort invested in creating software artifacts become obscured. Addressing these concerns is pivotal in maintaining a transparent and secure software development lifecycle.
Recently, there has been a growing research trend in detecting machine-generated texts [7, 8]. Perturbation-based methods like DetectGPT [9] have achieved state-of-the-art results in identifying machine-generated text. These methods employ likelihood score discrepancies between the original text and its various LLM-perturbed variants for detection, capturing the distinct patterns of machine-generated text that machines tend to prefer to a smaller set of phrases and expressions. However, such detection methods tailored for natural language texts face challenges when applied to code, as code requires strict adherence to syntactic rules, while natural language can maintain coherence with more variation [10]. This situation highlights a significant gap in existing research: a lack of in-depth assessment of the intrinsic features of machine- and human-authored code, crucial for understanding the unique patterns of machine-generated code and devising effective detection methods.
@@
In this paper, we conduct a comparative analysis of the distinct patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness. Through our analysis, we uncover that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens and adhere to common programming paradigms, and the disparity is more pronounced in stylistic tokens such as the whitespace tokens.
Based on the findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code. We extend the perturbation-based framework of DetectGPT by strategically inserting stylistic tokens, such as whitespace and newline characters, to capture the distinct patterns between machine- and human-authored code. This approach capitalizes on our observation that the disparity in coding styles is more pronounced in these stylistic tokens. By directly manipulating the code, DetectCodeGPT eliminates the need for an external pre-trained model, thereby enhancing both efficiency and effectiveness in the detection process.
To evaluate the effectiveness of DetectCodeGPT, we have conducted extensive experiments on two datasets across six code language models. The results demonstrate that DetectCodeGPT significantly outperforms the state-of-the-art methods by 7.6% in terms of AUC. Moreover, it proves to be a model-free and robust method against model discrepancies, making it viable for real-world applications with unknown or inaccessible source models.
@@
  • To our knowledge, we are the first to conduct a comprehensive and thorough analysis of the distinct patterns of LLM-generated code. Our study sheds light on essential insights that can further advance the utility of LLMs in programming.
  • We propose a novel method for detecting machine-generated code by leveraging its distinct stylistic patterns.
@@
II Background
II-A Large Language Models for Code
Large language models [11, 12] based on Transformer [13] decoder has achieved remarkable success in natural language processing tasks [14]. In the domain of code generation, Codex [1] and AlphaCode [15] are pioneering works to train large language models on code. The training data often contain millions of code in different programming languages collected from open source repositories like GitHub [16, 17, 18]. Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle [19, 20, 4, 5] and also instruction fine-tuning [21, 6]. Recent large language models pretrained on a mixture of programming and natural languages like ChatGPT [22] and LLaMA [23] have also shown promising results on code generation tasks.
II-B Perturbation-Based Detection of Machine-Generated Text
@@ III-A1 Lexical Diversity
We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human’s lexical usage.
III-A2 Conciseness
@@ III-A3 Naturalness
[…]
@@
III-B Experimental Setup
III-C Dataset Preparation
To compare with human-authored code, we extract 10,000 Python functions randomly from the CodeSearchNet corpus [16], which is curated from a wide range of open-source GitHub projects. We use the function signatures and their accompanying comments as prompts for the model as in [1]. We also collect the corresponding bodies of these functions to represent human-written code.
While acknowledging that current models, including CodeLlama and even ChatGPT [38], may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet [39, 38], the choice of this dataset is deliberate and insightful. Challenging the models against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [40] dataset, offers a more representative assessment. It allows us to analyze the differences between human- and machine-authored code when faced with broader, practical applications.
III-D Results and Analysis
We present the results and analysis regarding each code attribute introduced in Section III-A.
III-D1 Token Frequency
@@
Finding 1: Machine-authored code pays more attention to exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming paradigms.
III-D2 Syntax Element Distribution
Figure 1: Syntax element distribution of the code corpus
The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets. Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machine-authored code may have a more compact style with fewer identifiers.
Literal: Machine-authored code consistently shows a slightly higher tendency towards using literals. Across both temperatures, the machine code exhibits an increase in the literal proportion compared to the human-written code. This suggests that machine-authored code may be more likely to process raw data directly. This could result from the machine’s training on diverse data manipulation tasks.
@@ III-D2 Syntax Element Distribution
Finding 2: Machine-authored code tends to use fewer identifiers, more literals for direct data processing, and have more comments when the generation temperature grows.
III-D3 Zipf’s and Heaps’ Laws
@@
III-D4 Number of Tokens and Lines
@@
Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their stylistic preferences.
III-D5 Token Likelihood and Rank
Figure 4: Distribution of naturalness scores
@@ III-D5 Token Likelihood and Rank
IV Detecting Machine Generated Code
The empirical results suggest that machines tend to write more concise and natural code with a narrower spectrum of tokens, adhering to common programming paradigms, and the disparity is more pronounced in stylistic tokens such as whitespaces. This sparks a new idea for detecting machine-authored code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference. Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-authored code.
IV-A Problem Formulation
We formulate the detection of machine-authored code as a classification task, which predict whether or not a given code snippet $x$ is produced by a source model $p_{\theta}$. For this purpose, we transform $x$ to an equivalent form $\tilde{x}$ through a perturbation process $q(\cdot|x)$. We anticipate a sharper decline in its naturalness score if $x$ is written by an LLM. The key problems here are how to define the naturalness score and how to design the perturbation process. We introduce the naturalness score and the perturbation strategy $q(\cdot|x)$ in our approach in the following sections.
IV-B Measuring Naturalness
@@
IV-C Perturbation Strategy
Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style. Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below. Detailed explanations on the effectiveness of these perturbations are given in Section VI-A.
IV-C1 Space Insertion
$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (3)
IV-C2 Newline Insertion
We split the code into lines and obtain a set $L$ of lines. A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations. For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$. Formally, this is expressed as:
$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (4)
@@ 2       $p\sim\mathcal{U}(0,1)$;
4             Let $C$ represent all possible locations to insert spaces in $x$;
@@ 22      Store the perturbed code as $\tilde{x}_{i}$;
25Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
@@
We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations. Through this step, we instill randomness at a granular stylistic level, thereby amplifying the perturbation’s efficacy. Our perturbation strategy introduces several distinct advantages over the conventional methods [9, 41] using MLM to perturb the code, which will be discussed in Section VI-B.
@@ V Evaluation
  • RQ4: How effective is our method in cross-model code detection?
V-A Datasets
Apart from CodeSearchNet [16] described in Section III-C, we select Python code from The Stack [17] as another evaluation dataset. Similar to CodeSearchNet, The Stack provides code from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version [42] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1]. For each combination of dataset and model, we sample 500 human and machine code pairs for evaluation. The maximum length of code is trimmed to 128 tokens.
V-B Studied Models
We investigate machine-generated code by a diverse array of advanced LLMs, including Incoder [20], Phi-1 [43], StarCoder [44], WizardCoder [21], CodeGen2 [4] and CodeLlama [6]. We obtain their checkpoints from Huggingface  [^3] https://huggingface.co/models with different parameter sizes (1B-7B).
V-C Evaluation Metric
Following prior works [9, 41], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC). Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as: